OpenStudy (anonymous):

The amount of money a plumber charges when visiting a home to make a repair is shown by the linear equation c = 135 + 52h, where c is the amount the plumber charges and h is the number of hours spent at the home. How much does the plumber charge when h = 3.25? I WILL GIVE A MEDAL TO THE BEST ANSWER!!

OpenStudy (anonymous):

c=135 + 52(3.25) c=135 + 169 c=304